OpenStudy (anonymous):

Evaluate this expression. a = 4, b = -2, c = 8 ab - bc =

OpenStudy (anonymous):

You're given each variables value, so just put those in: 4(-2)+2(8)=?

OpenStudy (anonymous):

so -8

OpenStudy (anonymous):

It would be 8. 4*-2=-8 2*8=16 -8+16=8

OpenStudy (anonymous):

o srry i did it wrong thanks that makes sense now

OpenStudy (anonymous):

You're welcome